A particle moves on a straight line in such a way it covers 1st half distance with speed 3m/s and next half distance in two equal time intervals with speed 4.5m/s and 7.5m/s respectively. Find average speed of the particle.

Grade:11

1 Answers

Arun
6 years ago
 

let total distance of path is S ...

 

 

 

for first half , time taken if t

 

speed = distance/time

 

3 = S/2t

 

t = S/6      .............1

 

 

 

for second half , time taken is T ...

 

for first  T/2 it moves with 4.5

 

speed = distance/time

 

4.5 = 2x/T

 

 x = 4.5T/2   ...........2

 

for second T/2 it moves with 7.5

 

7.5 = 2y/T

 

y = 7.5T/2       ..........3

 

we have , x+y = S/2so

 

(7.5+4.5)T = S/2

 

T = S/24        ...............4

 

total time taken during its motion is T+t

 

 Ttotal = T+t = S/24+S/6 = 5S/24

 

total distance = S

 

average speed = total distance/total time = S/5S/24 = 24/5 = 4.8m/s